OpenStudy (anonymous):

log5 (x)= 4logx (5) , evaluate x.

OpenStudy (dls):

What do you mean? \[\Large \log_5 x=4 \log_x5\]?

OpenStudy (anonymous):

yup

OpenStudy (dls):

Do you know the base change formula?

OpenStudy (anonymous):

ya

OpenStudy (dls):

So can i write this like \[\Large \log_x5 =\frac{\log_55}{\log_5x}\]

OpenStudy (dls):

Do you know the value of \(\Huge \log_5 5\) ?

OpenStudy (anonymous):

1?

OpenStudy (dls):

yep,correct so now our equation reduces to : \[\Large \log_5 x=\frac{4}{\log_5x}\] Can you solve now?

OpenStudy (dls):

\[\Large (\log_5 x)^2=4\] Take sqrt on both sides etc can you solve?

OpenStudy (dls):

@kY_Tan

OpenStudy (anonymous):

log5 (x)= 2? but how to convert that log5?

OpenStudy (dls):

Nope, \[\Large \log_5x=\pm 2\]

OpenStudy (dls):

it can be -2 also...

OpenStudy (dls):

Now you gotta take antilog on both sides so raise both the sides to the power 5.

OpenStudy (dls):

\[\Huge 5^{\log_5x}=5^{2}\]

OpenStudy (anonymous):

every number after sqrt will be +and - together?

OpenStudy (dls):

yes mostly

OpenStudy (dls):

this is one solution,the other will be if you consider log(5)x=-2

OpenStudy (anonymous):

oh now i know , so the next step is log5 (x) = 2, right?

OpenStudy (anonymous):

@DLS

OpenStudy (dls):

i didn't get you? i did all the steps

OpenStudy (anonymous):

5^log5 (x) = 5^2 x=?

OpenStudy (dls):

you know on numerator you will get x and RHS is 25

OpenStudy (dls):

\[\Huge 5^{\log_5x}=x\]

OpenStudy (dls):

2nd solution you will get on equating with -2

OpenStudy (anonymous):

but why 25 is equal to x? the left hand side still have 5^log5(x) , sorry i don't understand this

OpenStudy (anonymous):

@DLS

OpenStudy (dls):

didn't you just said log base 5 (5) =1 similarly 5 to the power log base 5 (x)= x

OpenStudy (anonymous):

sorry, honestly i still don't understand how 5^log5 x is equal to x

OpenStudy (anonymous):

is it because 5 is an antilog then when it power a log 5 will be 1?

OpenStudy (dls):

its a property

OpenStudy (anonymous):

ok now i understood it , thanks @DLS

OpenStudy (dls):

yw :)

hero (hero):

Another way to do it: \(y = \log_5(x)\) is equivalent to \(x = 5^y\) \(y = 4\log_x(5)\) is equivalent to \(x^y = 5^4\) So we can work with \(x = 5^y\) and \(x^y = 5^4\) Since \(x = 5^y\) we substitute it in to the other equation to get: \((5^y)^y = 5^4\) \(5^{y^2} = 5^4\) So y = 2 Which means \(x = 5^2\)

OpenStudy (anonymous):

@Hero can i use this method to this question? "log2 (x)+ logx (4)=3"

hero (hero):

Not sure if it will work for this one.

OpenStudy (anonymous):

alright :)

OpenStudy (dls):

Is this the question? \[\Large \log_2 x + \log _x 4 =3 \] Please open a seperate one .

OpenStudy (mathmale):

And please, if at all possible, write the bases of logs properly: